OpenStudy (anonymous):

6x^2 + 7x -3? discriminant? i think the answer i got is wrong

OpenStudy (saifoo.khan):

121?

OpenStudy (anonymous):

no 120 :/

OpenStudy (saifoo.khan):

it's 121.

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

don't you mean 6^2

OpenStudy (anonymous):

so 36 + 84

OpenStudy (anonymous):

or...

OpenStudy (anonymous):

121

OpenStudy (anonymous):

how did you get that/?

OpenStudy (anonymous):

very hard this arithmetic. since when is \[7^2=36\]??

OpenStudy (saifoo.khan):

i just woke up actually. 3mins ago! =(

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

7^2 - 4(6)(-3) Solve..

OpenStudy (anonymous):

lol \[7^2-4\times 6\times (-3)\]

OpenStudy (anonymous):

Ok! wow thank you.

OpenStudy (anonymous):

the question is to find the discriminant. . . using the formula: \[b ^{2}-4ac\]to find the discriminant \[6x ^{2}+7x-3\] a=6 ; b=7; c=-3 --numerical terms of the quadratic equation given substitute the value solve: \[(7)^{2}-4(6)(-3)\] =121 ; final answer